# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ G.1 Proof of Lemma 2 : Implication of $(1-\delta$)-correctness
$\implies$ $\mathbb{P}(\boldsymbol{\pi}=\boldsymbol{\pi}_{\mathcal{F}}^{\star}\wedge\mathbf{A}\boldsymbol{\pi}_{\mathcal{F}}^{\star}\leq 0)\geq 1-\delta$
$\implies$ $\mathbb{P}(\mathbf{A}\boldsymbol{\pi}\leq 0)\geq 1-\delta$
@@ G.2 Stopping Criterion
$\vspace*{-.5em}\max_{\boldsymbol{\pi}_{t}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},\lambda_{a})>\beta(t,\delta)\vspace*{-.5em}$
where $\beta(t,\delta)\mathrel{\triangleq}3S_{0}\log(1+\log N_{a,t})+S_{0}\mathcal{T}\left(\frac{({K}\wedge d)+\log\frac{1}{\delta}}{S_{0}}\right)$, and $\rho(t,\delta)$ is in Lemma 4.
@@
$\boldsymbol{\pi}_{t}\mathrel{\triangleq}\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\hat{\mathcal{F}}}\hat{{\boldsymbol{\mu}}}_{t}^{\top}\boldsymbol{\pi}$, where $\hat{{\boldsymbol{\mu}}}_{t}\in\operatorname*{arg\,min}_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}N_{a,t}$.
@@
Algorithm 1 and 2 stops at a finite $\tau_{\delta}\in\mathbb{N}$ if the events $\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}(\hat{{\boldsymbol{\mu}}}_{t})}\{\exists t\in\mathbb{N}:\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})\}>\beta(t,\delta)$ and $\{\exists t\in\mathbb{N}:\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})N_{t}\|_{\infty}\leq t\rho(t,\delta)\}$ jointly occurs.
@@
Step 1: Finiteness of the stopping time. $\mathbf{A}$ stopping time is finite if the parameters in the system converges to their true values in finite time, in our case the means of arms and the constraint matrix. Let us define $\mathcal{\mathbf{A}}\mathrel{\triangleq}\{a\in[{K}]:\lim_{t\to\infty}N_{a,t}<\infty\}$ as a sampling rule i.e if an arm belongs to this set $\mathcal{\mathbf{A}}$, it has been sampled finitely and otherwise the arm has been sampled enough number of times so that the mean of that arm has converged to it’s true value and the column in the constraint matrix corresponding to that arm as also converged. For arms $a\in[{K}]\text{ and }a\in\mathcal{\mathbf{A}}^{c},\hat{{\boldsymbol{\mu}}}_{a,t}\rightarrow\tilde{{\boldsymbol{\mu}}}_{a}\neq{\boldsymbol{\mu}}_{a}$ and $(\tilde{\mathbf{A}})_{a,t}\to(\mathbf{A}^{\prime})_{a}\neq(\mathbf{A})_{a}$. When all parameters are concentrated $\mathcal{\mathbf{A}}=\emptyset$, we say $\forall a\in[{K}]:\hat{{\boldsymbol{\mu}}}_{a}\to{\boldsymbol{\mu}}_{a}$ and $\tilde{\mathbf{A}}\to\mathbf{A}$. We also define the limit of this empirical sampling rule as $\boldsymbol{\omega}_{\infty}=\lim_{t\to\infty}\frac{N_{a,t}}{t}\forall a\in[{K}]$. We then write the stopping condition in a new way $\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}(\hat{{\boldsymbol{\mu}}}_{t})}\{\exists t\in\mathbb{N}:\sum_{a=1}^{K}\frac{N_{a,t}}{t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>\frac{\beta(t,\delta)}{t}\}$ and $\{\exists t\in\mathbb{N}:\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\frac{N_{t}}{t}\|_{\infty}\leq\rho(t,\delta)\}$. By continuity properties and knowing $\beta(t,.)\to\log\log t$ and $\rho(t,.)\to 0$ as $t\to\infty$, we claim by taking taking asymptotic limits both sides $\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}\frac{\boldsymbol{\omega}_{\infty,a}}{t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>0$ and also $\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{\infty,a}\|_{\infty}<0$. We get strict inequality for the both the cases by the virtue of construction of the set $\mathcal{\mathbf{A}}$ such that for arms $a\in\mathcal{\mathbf{A}},\boldsymbol{\omega}_{\infty}\neq 0$ and the KL-divergence is non-zero as ${\boldsymbol{\lambda}}_{a}\neq{\boldsymbol{\mu}}$ since ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})$. Also for the second strict inequality, since $\tilde{\mathbf{A}}_{t}$ is the pessimistic estimate of $\mathbf{A}$ at time t the condition will hold.
@@
$\quad\qquad\qquad\qquad\wedge\left\{\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>\beta(t,\delta)\wedge\mathbf{A}\boldsymbol{\omega}_{t}\leq 0\right\}\Bigg{\}}$
@@
$\quad\qquad\qquad\qquad\wedge\left\{\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>\beta(t,\delta)\wedge\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{\infty}\leq\rho(t,\delta)\right\}\Bigg{\}}$
@@
$\mathbb{P}\{0\geq\mathbf{A}\boldsymbol{\omega}_{t}\}=\mathbb{P}\{-\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}\leq(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\}$$\leq\mathbb{P}\{\|\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}\|_{1}>\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{1}\geq\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{\infty}\}$
$\leq$ $\mathbb{P}\{\|\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}\|_{1}\leq\rho(t,\delta)\}\mathbb{P}\{\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{\infty}\leq\rho(t,\delta)\}$
=$=\mathbb{P}\{\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{\infty}\leq\rho(t,\delta)\}$
since the event $\{\|\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}\|_{1}\leq\rho(t,\delta)\}$ is a sure event.
@@
$\quad\mathbb{P}\left\{\exists t\in\mathbb{N}:\boldsymbol{\pi}_{t+1}=\boldsymbol{\pi}\wedge\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{\infty}>\rho(t,\delta)\right\}$
@@
$\quad+\sum_{\boldsymbol{\pi}\neq\boldsymbol{\pi}^{\star}_{\mathcal{F}}}\mathbb{P}\left\{\exists t\in\mathbb{N}:\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{\infty}>\rho(t,\delta)\right\}$ (17)
@@
$\sum_{\boldsymbol{\pi}\neq\boldsymbol{\pi}^{\star}_{\mathcal{F}}}\mathbb{P}\left\{\exists t\in\mathbb{N}:\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}_{t}\|_{\infty}>\rho(t,\delta)\right\}\leq\frac{1}{t}$
@@
since under ^$\hat{\mathcal{F}}_{t}$ and the bad event we are assuming that the estimated alt-set is still bigger than the actual alt-set. So any ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})$ will have a bigger distance with the estimate of ${\boldsymbol{\mu}}$ than any ${\boldsymbol{\lambda}}^{\prime}\in\Lambda_{\mathcal{F}}(\hat{{\boldsymbol{\mu}}}_{t})$. We define $I_{\boldsymbol{\pi}}\mathrel{\triangleq}\mathrm{Supp}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})\Delta\mathrm{Supp}(\boldsymbol{\pi})$ and also $S_{0}\mathrel{\triangleq}\max_{\boldsymbol{\pi}}|I_{\boldsymbol{\pi}}|$. We note that $0\leq S_{0}\leq{K}$.
@@ G.3 Upper Bound of LATS
where $\mathfrak{s}$ is the shadow price of the true constraint $\mathbf{A}$, $T_{\mathcal{F},r}(\mu)$ is the characteristic time under the true constraint (Equation (3.2)), and $\delta\in(0,1]$.
@@
$G_{T}\triangleq\bigcap_{t=h(T)}^{T}\{||\hat{{\boldsymbol{\mu}}}_{t}-{\boldsymbol{\mu}}||_{\infty}\leq\xi(\epsilon)\wedge||(\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}||_{\infty}\leq\phi(\epsilon)\forall\boldsymbol{\omega}\in\hat{\mathcal{F}}\}$
@@
We also observe that $||{\boldsymbol{\mu}}^{\prime}-{\boldsymbol{\mu}}||_{\infty}\leq\xi(\epsilon)$ and $||(\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}||_{\infty}\leq\phi(\epsilon)$ implies that $\sup\limits_{\boldsymbol{\omega}^{\prime}\in\boldsymbol{\omega}^{\star}({\boldsymbol{\mu}}^{\prime})}\sup\limits_{\boldsymbol{\omega}\in\boldsymbol{\omega}^{\star}({\boldsymbol{\mu}})}||\boldsymbol{\omega}^{\prime}-\boldsymbol{\omega}||\leq\epsilon$ due to upper hemicontinuity of $\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$ (Theorem 3).
@@
$\mathbb{P}(G_{T}^{c})$ $=\sum_{t=h(T)}^{T}\left(\mathbb{P}\left\{||\hat{{\boldsymbol{\mu}}}_{t}-{\boldsymbol{\mu}}||_{\infty}>\xi(\epsilon)\right\}+\mathbb{P}\left\{||(\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}||_{\infty}>\phi(\epsilon)\right\}\right)$
$\leq\sum_{t=h(T)}^{T}\mathbb{P}\left\{||\hat{{\boldsymbol{\mu}}}_{t}-{\boldsymbol{\mu}}||_{\infty}>\xi(\epsilon)\right\}+\sum_{t=h(T)}^{T}\mathbb{P}\left\{||(\tilde{\mathbf{A}}_{t}^{i}-\mathbf{A}^{i})\boldsymbol{\omega}||_{\infty}>\phi(\epsilon)\right\}$
$\leq BT\exp\left(-CT^{\frac{1}{8}}\right)+{K}\sum_{t=h(T)}^{T}\frac{1}{t}$
The first inequality is due to the union bound. The second inequality is due to the Lemma  7 (Lemma 19 of Garivier and Kaufmann, (2016)), which states that
$\sum_{t=h(T)}^{T}\mathbb{P}\left\{\|\hat{{\boldsymbol{\mu}}}_{t}-{\boldsymbol{\mu}}\|_{\infty}>\xi(\epsilon)\right\}\leq BT\exp\left(-CT^{\frac{1}{8}}\right)\,,$
@@
Step 3: Tracking argument. Now, we state how concentrating on means and constraints leads to good concentration on the allocations too. Since we use C-tracking, we can leverage the concentration in allocation by (Degenne et al., 2019b, , Lemma 17). We use this lemma than D-tracking or the tracking argument in  (Garivier and Kaufmann,, 2016, Lemma 7) because the optimal allocations might not be unique but the set $\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$ is convex (Theorem 3).
@@
We have replaced the perturbed alt-set with the true one because for $t\geq\max(T_{\epsilon},h(T))$, we can ensure the convergence of $\hat{\mathcal{F}}$ almost surely i.e ^$\hat{\mathcal{F}}\overset{\mathrm{a.s}}{\to}\mathcal{F}$
@@
$C_{\epsilon,\hat{\mathcal{F}}}\mathrel{\triangleq}\inf_{\begin{subarray}{c}{\boldsymbol{\mu}}^{\prime}:||{\boldsymbol{\mu}}^{\prime}-{\boldsymbol{\mu}}||_{\infty}\leq\xi(\epsilon)\\ \boldsymbol{\omega}^{\prime}:||\boldsymbol{\omega}^{\prime}-\boldsymbol{\omega}||_{\infty}\leq 3\epsilon\\ \tilde{\mathbf{A}}^{\prime}:||(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}||_{\infty}\leq\phi(\epsilon)\end{subarray}}\mathcal{D}({\boldsymbol{\mu}}^{\prime},\boldsymbol{\omega}^{\prime},\hat{\mathcal{F}})-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\,.$
@@
$(1+\psi)\inf_{\begin{subarray}{c}{\boldsymbol{\mu}}^{\prime}:||{\boldsymbol{\mu}}^{\prime}-{\boldsymbol{\mu}}||_{\infty}\leq\xi(\epsilon)\\ \boldsymbol{\omega}^{\prime}:||\boldsymbol{\omega}^{\prime}-\boldsymbol{\omega}||_{\infty}\leq 3\epsilon\\ \tilde{\mathbf{A}}:||(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}||_{\infty}\leq\phi(\epsilon)\end{subarray}}\mathcal{D}({\boldsymbol{\mu}}^{\prime},\boldsymbol{\omega}^{\prime},\hat{\mathcal{F}})\geq\inf_{\begin{subarray}{c}{\boldsymbol{\mu}}^{\prime}:||{\boldsymbol{\mu}}^{\prime}-{\boldsymbol{\mu}}||_{\infty}\leq\xi(\epsilon)\\ \boldsymbol{\omega}^{\prime}:||\boldsymbol{\omega}^{\prime}-\boldsymbol{\omega}||_{\infty}\leq 3\epsilon\\ \tilde{\mathbf{A}}^{\prime}:||(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}||_{\infty}\leq\phi(\epsilon)\end{subarray}}\mathcal{D}({\boldsymbol{\mu}}^{\prime},\boldsymbol{\omega}^{\prime},\hat{\mathcal{F}})-\boldsymbol{l}^{\top}\tilde{\mathbf{A}}\boldsymbol{\omega}\,,$
@@
$\mathcal{D}(\hat{{\boldsymbol{\mu}}}_{t},N_{t},\hat{\mathcal{F}})\geq t\inf_{\begin{subarray}{c}{\boldsymbol{\mu}}^{\prime}:||{\boldsymbol{\mu}}^{\prime}-{\boldsymbol{\mu}}||_{\infty}\leq\xi(\epsilon)\\ \boldsymbol{\omega}^{\prime}:||\boldsymbol{\omega}^{\prime}-\boldsymbol{\omega}||_{\infty}\leq 3\epsilon\\ \tilde{\mathbf{A}}:||(\tilde{\mathbf{A}}-\mathbf{A})\boldsymbol{\omega}||_{\infty}\leq\phi(\epsilon)\end{subarray}}\mathcal{D}({\boldsymbol{\mu}}^{\prime},\boldsymbol{\omega}^{\prime},\hat{\mathcal{F}})\geq\frac{tC_{\epsilon,\hat{\mathcal{F}}}}{1+\psi}\,.$ (18)
@@
$\min(\tau_{\delta},T)\leq\max(\sqrt{T},f(T,\delta))+\sum_{t=T_{\epsilon}}^{T}\mathds{1}_{\tau_{\delta}>t}$
@@
Now using the lower bound on $\mathcal{D}(\hat{{\boldsymbol{\mu}}}_{t},N_{t},\hat{\mathcal{F}})$ (Equation 18), we get
$\min(\tau_{\delta},T)\leq\beta(t,\delta))$ $\leq\max(\sqrt{T},\beta(T,\delta))+\sum_{t=T_{\epsilon}}^{T}\mathds{1}\Bigl{(}t\frac{C_{\epsilon,\hat{\mathcal{F}}}}{1+\psi}\leq\beta(t,\delta)\Bigr{)}$
$\leq\max(\sqrt{T},\beta(T,\delta))+\sum_{t=T_{\epsilon}}^{T}\mathds{1}\Bigl{(}t\frac{C_{\epsilon,\hat{\mathcal{F}}}}{1+\psi}\leq\beta(T,\delta)\Bigr{)}$
@@
Now, if also $\epsilon\rightarrow 0$, by the Equation  4, we get $\tilde{\mathbf{A}}\rightarrow\mathbf{A}$, and thus, ^$\hat{\mathcal{F}}\rightarrow F$.

### v2

@@ G.1 Proof of Lemma 2 : Implication of $(1-\delta)$-correctness
$\implies$ $\mathbb{P}(\boldsymbol{\pi}=\boldsymbol{\pi}_{\mathcal{F}}^{\star}\wedge A\boldsymbol{\pi}_{\mathcal{F}}^{\star}\leq 0)\geq 1-\delta$
$\implies$ $\mathbb{P}(A\boldsymbol{\pi}\leq 0)\geq 1-\delta$
@@ G.2 Stopping Criterion
[…]
where $\beta(t,\delta)\mathrel{\triangleq}3S_{0}\log(1+\log N_{a,t})+S_{0}\mathcal{T}\left(\frac{({K}\wedge d)+\log\frac{1}{\delta}}{S_{0}}\right)$ and $0\leq S_{0}\leq{K}$.
@@
$\boldsymbol{\pi}_{t}\mathrel{\triangleq}\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\hat{\mathcal{F}}}\hat{{\boldsymbol{\mu}}}_{t}^{\top}\boldsymbol{\pi}$, where $\hat{{\boldsymbol{\mu}}}_{t}\in\operatorname*{arg\,min}_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}N_{a,t}$.
@@
Algorithm 2 and 3 stops at a finite $\tau_{\delta}\in\mathbb{N}$ if the events $\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}(\hat{{\boldsymbol{\mu}}}_{t})}\{\exists t\in\mathbb{N}:\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})\}>\beta(t,\delta)$ and $\{\exists t\in\mathbb{N}:\|(\tilde{A}_{t}-A)N_{t}\|_{\infty}\leq t\rho(t,\delta)\}$ jointly occurs.
@@
Step 1: Finiteness of the stopping time. $A$ stopping time is finite if the parameters in the system converges to their true values in finite time, in our case the means of arms and the constraint matrix. Let us define $\mathcal{A}\mathrel{\triangleq}\{a\in[{K}]:\lim_{t\to\infty}N_{a,t}<\infty\}$ as a sampling rule i.e if an arm belongs to this set $\mathcal{A}$, it has been sampled finitely and otherwise the arm has been sampled enough number of times so that the mean of that arm has converged to it’s true value and the column in the constraint matrix corresponding to that arm as also converged. For arms $a\in[{K}]\text{ and }a\in\mathcal{A}^{c},\hat{{\boldsymbol{\mu}}}_{a,t}\rightarrow\tilde{{\boldsymbol{\mu}}}_{a}\neq{\boldsymbol{\mu}}_{a}$ and $(\tilde{A})_{a,t}\to(A^{\prime})_{a}\neq(A)_{a}$. When all parameters are concentrated $\mathcal{A}=\emptyset$, we say $\forall a\in[{K}]:\hat{{\boldsymbol{\mu}}}_{a}\to{\boldsymbol{\mu}}_{a}$ and $\tilde{A}\to A$. We also define the limit of this empirical sampling rule as $\boldsymbol{\omega}_{\infty}=\lim_{t\to\infty}\frac{N_{a,t}}{t}\forall a\in[{K}]$. We then write the stopping condition in a new way $\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}(\hat{{\boldsymbol{\mu}}}_{t})}\{\exists t\in\mathbb{N}:\sum_{a=1}^{K}\frac{N_{a,t}}{t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>\frac{\beta(t,\delta)}{t}\}$ and $\{\exists t\in\mathbb{N}:\|(\tilde{A}_{t}-A)\frac{N_{t}}{t}\|_{\infty}\leq\rho(t,\delta)\}$. By continuity properties and knowing $\beta(t,.)\to\log\log t$ and $\rho(t,.)\to 0$ as $t\to\infty$, we claim by taking taking asymptotic limits both sides $\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}\frac{\boldsymbol{\omega}_{\infty,a}}{t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>0$ and also $\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{\infty,a}\|_{\infty}<0$. We get strict inequality for the both the cases by the virtue of construction of the set $\mathcal{A}$ such that for arms $a\in\mathcal{A},\boldsymbol{\omega}_{\infty}\neq 0$ and the KL-divergence is non-zero as ${\boldsymbol{\lambda}}_{a}\neq{\boldsymbol{\mu}}$ since ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})$. Also for the second strict inequality, since $\tilde{A}_{t}$ is the optimistic estimate of $A$ at time t the condition will hold.
@@
$\quad\qquad\qquad\qquad\wedge\left\{\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>\beta(t,\delta)\wedge A\boldsymbol{\omega}_{t}\leq 0\right\}\Bigg\}$
@@
$\quad\qquad\qquad\qquad\wedge\left\{\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})>\beta(t,\delta)\wedge\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{\infty}\leq\rho(t,\delta)\right\}\Bigg\}$
@@
$\mathbb{P}\{0\geq A\boldsymbol{\omega}_{t}\}=\mathbb{P}\{-\tilde{A}_{t}\boldsymbol{\omega}_{t}\leq(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\}$ $\leq\mathbb{P}\{\|\tilde{A}_{t}\boldsymbol{\omega}_{t}\|_{1}>\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{1}\geq\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{\infty}\}$
$\leq$ $\mathbb{P}\{\|\tilde{A}_{t}\boldsymbol{\omega}_{t}\|_{1}\leq\rho(t,\delta)\}\mathbb{P}\{\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{\infty}\leq\rho(t,\delta)\}$
$=\mathbb{P}\{\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{\infty}\leq\rho(t,\delta)\}$
since the event $\{\|\tilde{A}_{t}\boldsymbol{\omega}_{t}\|_{1}\leq\rho(t,\delta)\}$ is a sure event.
@@
$\quad\mathbb{P}\left\{\exists t\in\mathbb{N}:\boldsymbol{\pi}_{t+1}=\boldsymbol{\pi}\wedge\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{\infty}>\rho(t,\delta)\right\}$
@@
$\quad+\sum_{\boldsymbol{\pi}\neq\boldsymbol{\pi}^{\star}_{\mathcal{F}}}\mathbb{P}\left\{\exists t\in\mathbb{N}:\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{\infty}>\rho(t,\delta)\right\}$ (15)
@@
$\sum_{\boldsymbol{\pi}\neq\boldsymbol{\pi}^{\star}_{\mathcal{F}}}\mathbb{P}\left\{\exists t\in\mathbb{N}:\|(\tilde{A}_{t}-A)\boldsymbol{\omega}_{t}\|_{\infty}>\rho(t,\delta)\right\}\leq\frac{1}{t}$
@@
since under $\hat{\mathcal{F}}_{t}$ and the bad event we are assuming that the estimated alt-set is still bigger than the actual alt-set. So any ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})$ will have a bigger distance with the estimate of $\boldsymbol{\mu}$} than any ${\boldsymbol{\lambda}}^{\prime}\in\Lambda_{\mathcal{F}}(\hat{{\boldsymbol{\mu}}}_{t})$. We define $I_{\boldsymbol{\pi}}\mathrel{\triangleq}\mathrm{Supp}(\boldsymbol{\pi}^{\star}_{\mathcal{F}})\Delta\mathrm{Supp}(\boldsymbol{\pi})$ and also $S_{0}\mathrel{\triangleq}\max_{\boldsymbol{\pi}}|I_{\boldsymbol{\pi}}|$. We note that $0\leq S_{0}\leq{K}$.
@@ G.3 Upper Bound of LATS
where $\mathfrak{s}$ is the shadow price of the true constraint $A$, $T_{\mathcal{F},r}(\mu)$ is the characteristic time under the true constraint (Equation (4)), and $\delta\in(0,1]$.
@@
[…]
@@
[…]
@@
[…]
$\leq BT\exp\left(-CT^{\frac{1}{8}}\right)+{K}\sum_{t=h(T)}^{\top}\frac{1}{t}$
The first inequality is due to the union bound. The second inequality is due to the Lemma 7 (Lemma 19 of Garivier and Kaufmann, (2016)), which states that
$\sum_{t=h(T)}^{\top}\mathbb{P}\left\{\|\hat{{\boldsymbol{\mu}}}_{t}-{\boldsymbol{\mu}}\|_{\infty}>\xi(\epsilon)\right\}\leq BT\exp\left(-CT^{\frac{1}{8}}\right)\,$,
@@
Step 3: Tracking argument. Now, we state how concentrating on means and constraints leads to good concentration on the allocations too. Since we use C-tracking, we can leverage the concentration in allocation by(Lemma 17 of Degenne et al., 2019b ). We use this lemma than D-tracking or the tracking argument in (Lemma 7 of Garivier and Kaufmann, (2016)) because the optimal allocations might not be unique but the set $\boldsymbol{\omega}^{*}({\boldsymbol{\mu}})$ is convex (Theorem 3).
@@
We have replaced the perturbed alt-set with the true one because for $t\geq\max(T_{\epsilon},h(T))$, we can ensure the convergence of $\hat{\mathcal{F}}$ almost surely i.e $\hat{\mathcal{F}}\overset{\mathrm{a.s}}{\to}\mathcal{F}$
@@
[…]
@@
[…]
@@
[…]
@@
$\min(\tau_{\delta},T)\leq\max(\sqrt{T},f(T,\delta))+\sum_{t=T_{\epsilon}}^{\top}\mathds{1}_{\tau_{\delta}>t}$
@@
[…]
$\min(\tau_{\delta},T)\leq\beta(t,\delta))$ $\leq\max(\sqrt{T},\beta(T,\delta))+\sum_{t=T_{\epsilon}}^{\top}\mathds{1}\Bigl(t\frac{C_{\epsilon,\hat{\mathcal{F}}}}{1+\psi}\leq\beta(t,\delta)\Bigr)$
$\leq\max(\sqrt{T},\beta(T,\delta))+\sum_{t=T_{\epsilon}}^{\top}\mathds{1}\Bigl(t\frac{C_{\epsilon,\hat{\mathcal{F}}}}{1+\psi}\leq\beta(T,\delta)\Bigr)$
@@
Now, if also $\epsilon\rightarrow 0$, by the Equation 3, we get $\tilde{A}\rightarrow A$, and thus, $\hat{\mathcal{F}}\rightarrow F$.